### v1

# CLEAR-IR: Clarity-Enhanced Active Reconstruction of Infrared Imagery

Abstract
This paper presents a novel approach for enabling robust robotic perception in dark environments using infrared (IR) stream. IR stream is less susceptible to noise than RGB in low-light conditions. However, it is dominated by active emitter patterns that hinder high-level tasks such as object detection, tracking and localisation. To address this, a U-Net-based architecture is proposed that reconstructs clean IR images from emitter-populated input, improving both image quality and downstream robotic performance. This approach outperforms existing enhancement techniques and enables reliable operation of vision-driven robotic systems across illumination conditions from well-lit to extreme low-light scenes.
I Introduction
Lighting-invariant vision systems are desirable for enabling robots to operate robustly across diverse and unpredictable environments without requiring modifications to the underlying perception pipeline. In order to support high-level tasks such as object detection, semantic segmentation, and image classification, the vision system must remain reliable even in low light or completely dark scenes. Such capabilities are critical in domains like mine shaft exploration, post-disaster victim identification, nuclear facility inspection, and visual loop closure in feature-deprived environments using aruco markers.
Conventional RGB cameras, while effective under well-lit conditions, suffer from significant degradation in low-light scenes [1]. Image quality is compromised due to sensor-level limitations, such as dark current and readout noise. Attempts to compensate using an increased exposure time introduce motion blur, rendering them ineffective for dynamic environments. Furthermore, data-driven approaches that require training networks in long-exposure or low-light image pairs add complexity and reduce generalisability across different camera models. In addition, such methods are fundamentally limited by information-theoretic constraints and if critical scene information is not captured by the sensor due to noise or extreme darkness, no amount of training data can fully reconstruct it. The effectiveness of these networks is therefore heavily dependent on both data availability and inherent scene observability.
Another approach involves equipping robots with onboard illumination systems [2]. However, in certain rich environments such as underground tunnels or disaster zones, the Tyndall effect causes the light to scatter off airborne particles, saturating the camera view with undesirable visual artefacts and occluding key features in the scene. This method also suffers from overexposure, specular reflection, and increased power consumption. Moreover, the use of active illumination can produce strong shadows that obscure structural details and complicate downstream vision tasks such as object detection and localisation. To address these challenges, this work explores the use of infrared (IR) imaging, particularly active IR systems that operate independently of ambient illumination. As seen in Figure 1, the dark scene captured by the RGB camera fails to present good visual cues, whereas the same scene in the IR spectrum can be visualised with several objects in the scene. Although IR cameras are promising for dark-scene perception, they are not without their own limitations. Passive IR sensors, which operate without emitters, avoid structured-light artefacts but are typically expensive and limited to low-resolution outputs, making them impractical for robotics. Active IR systems, introduce structured light patterns that degrade image quality, interfere with feature extraction, and disrupt high-level perception tasks such as object detection, tracking, and SLAM.
To overcome this, a deep learning-based framework, CLEAR-IR (Clarity-Enhanced Active Reconstruction of Infrared Imagery), is proposed. The framework is designed to reconstruct clean and interpretable IR images from raw emitter-populated inputs. The method leverages a U-Net architecture trained to suppress structured light noise while preserving spatial and textural details critical for vision-based tasks. The output is a denoised, visually coherent IR image that can serve as a robust input to standard perception modules. This is the first work to leverage the infrared stream for high-level robotic tasks by removing active emitter patterns in scenes ranging from well-illuminated to extreme low-light. The contributions of this work are:
  • CLEAR-IR, a U-Net–based framework tailored for reconstructing clean IR images by suppressing structured-light artefacts, is introduced.
  • A composite loss function is developed to preserve spatial and textural details critical for robotic perception while robustly removing high-frequency emitter patterns.
  • Evaluation of IR image reconstruction for robotics is presented, demonstrating significant improvements across key tasks such as feature matching, object detection, and marker detection.
(a) Low-light room captured using an RGB camera
(b) Low-light room captured using an IR camera
Figure 1: Low-light room captured using different cameras
II Related Work
Vision-based robotic perception in challenging lighting conditions has been an active area of research. One line of work has focused on improving the performance of standard RGB cameras in low light. This includes sensor-level enhancements such as improved read-out noise reduction and pixel architectures [3], as well as computational photography techniques such as multiexposure fusion [4] and burst photography [5]. Although these techniques can improve image quality, they are often computationally intensive and can introduce motion blur or require static scenes, limiting their use in dynamic robotic applications.
A separate body of work leverages deep learning for low-light image enhancement. Prominent examples include supervised learning methods that train neural networks on paired low light and well lit images [6, 7]. These approaches, however, are highly dependent on the quality and diversity of the training data and may not generalise well to different camera hardware or unseen environments. Unsupervised and semi-supervised methods have also been explored to mitigate the need for paired data as seen in [8, 9, 10], but they often struggle to remove complex artefacts such as structured light patterns found in active IR systems, which this work explores. Generative adversarial networks (GANs) have been used for enhancement in low light as inferred from [11, 12, 13], but these methods do not translate well to extremely low light scenes, which require some form of active sensing. Moreover, the artefacts introduced by these methods are particularly problematic for robotics: vision-based SLAM pipelines may incorrectly latch onto these artificial structures as features, leading to spurious correspondences, tracking drift, and eventual map corruption. Thus, preserving fine-grained but genuine details is critical for ensuring robustness in downstream tasks such as vSLAM and visual odometry.
Active infrared (IR) systems, which operate by illuminating a scene with an IR light source and capturing the reflected light, have been a viable alternative for dark-scene perception. These systems have been widely adopted in applications such as facial recognition [14] and depth sensing [15]. However, the structured light patterns they use for scene illumination and depth estimation, while beneficial for some tasks, present a significant challenge for 2D vision tasks like object detection and segmentation. Repetitive high-frequency patterns can be misinterpreted as texture or objects by standard vision algorithms, leading to false positives or missed detections, which limits the amount of work done using these sensors. Previous works have only attempted to remove fixed patterns from IR images, as seen in [16, 17, 18]. This just enhances the extremely noisy input of IR images, but does not focus on making it more useful for high-level robotics tasks.
The U-Net architecture, with its symmetric encoder-decoder structure and skip connections, has shown remarkable success in a variety of image-to-image translation tasks, and was first introduced for biomedical image segmentation in [19]. Its ability to capture both high-level semantic information from the encoder path and fine-grained spatial details from the decoder path makes it an ideal candidate for the specific denoising problem encountered in IR images. This architecture has also been successfully adapted for other image reconstruction tasks, such as super-resolution [20] and general image denoising [21].
In contrast to existing methods that primarily focus on low-light enhancement or general noise removal, this work addresses the specific and unique problem of structured light artefacts in active IR imagery. Using the U-Net architecture, it is possible to effectively learn to suppress these specific patterns while preserving the crucial underlying semantic information for high-level vision tasks.
III Methodology
This section presents the CLEAR-IR framework. This method leverages a U-Net architecture and a composite loss function to effectively remove structured light artefacts while preserving critical visual details.
III-A Emitter Pattern Characteristics
Active infrared (IR) systems enable vision in dark environments by projecting an illumination pattern onto the scene. This pattern is often a pseudo-random pattern which is a semi-deterministic and repeatable sequence of dots generated by an algorithm. The raw IR image captured by the camera is, therefore, populated with these dots, which can be misidentified as texture or objects by standard vision algorithms, complicating tasks like object detection and segmentation.
The intensity of these dots in the captured image is not uniform and depends on four main factors:
  • Object Reflectivity: Objects with higher reflectivity in the IR spectrum will appear brighter, as they reflect a sharper emitter pattern to the camera, as shown in Figure2a, where the plastic screen reflects more emitter pattern than the black foam.
  • Distance to the Emitter: The intensity of the light from the emitter decreases with the square of the distance, following the inverse square law. Consequently, closer objects will show a more intense and prominent dot pattern, where the patterns are larger when closer and smaller when they move further away from the camera, as seen in Figure2b and Figure2a, where the closer part of the sheet has more shaper patterns than the further part.
  • Emitter Power: The overall intensity of the pattern is directly proportional to the laser power of the emitter, where a higher power leads to a denser, more defined pattern. This is also a critical factor in the appearance of the pattern, as seen in Figure 2c, where a lower laser power prevents the room from being seen while a larger laser power populates the room with the emitter pattern
  • Room Brightness: This factor contributes to the emitter power configuration, where the same emitter pattern may appear dimmer in well-illuminated scenes while appearing sharper in low-light scenes, as seen in Figure2d and Figure2e, where ambient light affects the emitter visibility by dimming the laser power.
(a) Emitter pattern’s appearance based on object reflectivity
(b) Emitter pattern’s appearance based on distance
(c) Changes in the emitter pattern’s intensity based on the laser power of the emitter
(d) Emitter pattern in a bright room
(e) Emitter pattern in a dark room
Figure 2: The key factors influencing the appearance of the structured light emitter pattern in active infrared imagery.
This variability in pattern appearance across different scenes creates a significant challenge in selecting adequate losses that can remove them from the images while being scene-agnostic.
III-B U-Net Architecture for Emitter Pattern Removal
To effectively reconstruct the image and remove the noise caused by the emitter pattern, a standard U-Net architecture was used. This convolutional neural network is particularly well suited for image-to-image translation tasks due to its symmetric encoder-decoder structure and strategic use of skip connections, as depicted in Figure 3.
The network is configured to process greyscale images with an input resolution of 640x480 pixels. The encoder path progressively downsamples the image, using a series of convolutional and max-pooling layers to extract increasingly abstract features and high-level contextual information.
The decoder path then reconstructs the image using transposed convolutional layers to upsample the features back to the original resolution. A crucial element of this architecture are its skip connections, which concatenate the feature maps from the encoder to the corresponding layers in the decoder. These connections are vital for preserving fine-grained spatial details that might otherwise be lost during the downsampling process. The network combines these features at each upsampling stage to produce a final, single-channel output image that is the same size as the input, with the emitter pattern effectively removed.
Figure 3: The U-Net architecture used for image reconstruction and emitter pattern removal.
III-C Dataset and Data Augmentation
A custom paired dataset of IR images and their corresponding clean ground-truth RGB images converted to greyscale was created for model training using an Intel Realsense D455 camera [22]. Data collection was performed in an indoor workspace with varied lighting conditions. The illumination levels and object distance were intentionally adjusted to capture emitter patterns with different sharpness and intensity, creating a varied dataset.
The initial dataset contained 10,000 paired images. To enhance model robustness and prevent overfitting, a data augmentation process was applied. For each image pair, five augmented versions were generated using random transformations, which included:
  • Rotation: Up to 30 degrees in either direction.
  • Flipping: Random horizontal and vertical flips.
  • Affine Transformations: Minor rotations and translations.
  • Zoom: Random zooming in by up to 30%.
  • Brightness and Contrast: Adjustments to simulate lighting variations.
The final data had 50000 paired sets of IR and greyscale RGB images and will be released as an open-sourced dataset.
III-D Composite Loss Function and Training
To produce high-quality image reconstructions, a composite loss function was used to guide the network’s training. This function is a weighted sum of multiple components, each designed to optimise a different aspect of the output, ensuring that the reconstructed image is not pixel accurate but perceptually and structurally similar to the ground truth. The main reason for choosing these losses besides mean absolute or mean squared error was due to the epipolar transformation of pixels between the IR and RGB images.
III-D1 Loss Components
The total loss is a weighted sum of several components. Let $\alpha,\beta,\gamma,\delta,\epsilon$, and $\zeta$ be the empirically tuned weighting coefficients for each corresponding loss component. The total loss is defined as:
[…]
  • Mean Absolute Error (MAE): This is a pixel-wise loss that minimises the average absolute difference between the predicted image $\hat{I}$ and the ground truth image $I$. It provides a foundational measure of pixel level accuracy.
    $\mathcal{L}_{mae}=\frac{1}{N}\sum_{i=1}^{N}|\hat{I}_{i}-I_{i}|$ (4)
  • Structural Similarity Index (SSIM) Loss: This loss measures the similarity in luminance, contrast, and structure between images. By minimising this loss, the network is encouraged to preserve the overall structural integrity of the scene [23].
    $\mathcal{L}_{ssim}=1-\text{SSIM}(\hat{I},I)$ (5)
  • High-Frequency and Sobel Edge Losses: These losses are used to preserve fine details and sharp boundaries. The high-frequency loss uses a Laplacian filter to penalise differences in the high-frequency components of the images, while the Sobel edge loss penalises differences in image gradients to ensure edges are accurately reproduced.
    $\mathcal{L}_{freq}=\left\|(\hat{I}-I)*K_{Laplacian}\right\|_{1}$ (6)
    $\mathcal{L}_{sobel}=\left\|\nabla_{x}\hat{I}-\nabla_{x}I\right\|_{1}+\left\|\nabla_{y}\hat{I}-\nabla_{y}I\right\|_{1}$ (7)
  • Perceptual Loss: This loss component uses a pre-trained VGG19 network as a feature extractor. The network’s weights, which were pre-trained on the ImageNet dataset [24], are frozen during training. This allows the model to leverage learned feature representations that are more aligned with human visual perception than simple pixel values. To apply this to greyscale images, the input is converted to a 3 channel format by concatenating the single channel three times. Instead of comparing pixels directly, this loss compares the feature maps from a set of specific intermediate layers within the VGG19 network. By using layers such as block3_conv2, block4_conv2, and the layers within block5, the loss encourages the reconstructed image to have similar high level content and style as the ground truth. This is crucial for generating an output that is visually convincing and suitable for high level vision tasks [25].
    $\mathcal{L}_{perceptual}=\sum_{j}\|\Phi_{j}(\hat{I})-\Phi_{j}(I)\|_{1}$ (8)
    where $\Phi_{j}(\cdot)$ is the feature map from layer $j$ of the VGG19 network.
  • Total Variation (TV) Loss: This regularizer promotes spatial smoothness and helps to reduce spurious noise in the output image [26].
    $\mathcal{L}_{tv}=\sum_{i,j}\left(\left(\hat{I}_{i+1,j}-\hat{I}_{i,j}\right)^{2}+\left(\hat{I}_{i,j+1}-\hat{I}_{i,j}\right)^{2}\right)$ (9)
III-D2 Weighting Strategy
The weighting coefficients for each component ($\alpha,\beta,\gamma,\delta,\epsilon,\zeta$) were experimentally tuned to balance their influence on the final output. The primary goal was to prioritise visual quality and structural integrity over a strict pixel-by-pixel match. High weights were assigned to $\mathcal{L}_{ssim}$ to preserve scene structure, while a significant but not overwhelming weight was given to $\mathcal{L}_{perceptual}$ to ensure the output was visually convincing. Moderate weights were used for $\mathcal{L}_{freq}$ and $\mathcal{L}_{sobel}$ to preserve fine details and edges while also not prioritising emitter patterns which can be classed as high-frequency patterns in the image. In contrast, very low weights were given to $\mathcal{L}_{mae}$ and $\mathcal{L}_{tv}$, as their purpose was to act as subtle regularizers rather than primary training objectives.
III-E Preprocessing and Training
After augmentation, the dataset, which contained 50,000 paired images, 90% of the data was split into a training set and the remaining 10% into the validation set. All images were resized to a consistent resolution of 640x480 pixels and normalised to the range of [0, 1].
The model was trained using the Adam optimizer with a CosineDecayRestarts learning rate schedule to minimise the composite loss function. To prevent overfitting, an early stopping mechanism with a patience of 30 epochs was implemented.
The model was trained for a total of 126 epochs, at which point it met the early stopping criteria, indicating that a sufficient level of convergence had been reached.
IV Experimental Results
This section presents a comprehensive evaluation of the proposed CLEAR-IR framework, focusing on its effectiveness in preparing infrared images for high-level computer vision tasks. Since this is the first work to specifically address the removal of active emitter patterns, the primary evaluation is a qualitative and quantitative comparison of the performance of enhancement techniques on the RGB stream versus the denoised images of the IR stream produced by this work.
IV-A Qualitative and Quantitative Analysis
To provide a visual assessment of the method’s reconstruction quality and its robustness in low light conditions, a qualitative comparison is first presented of a raw IR input, the model’s reconstructed output, and the corresponding ground truth image, as seen in Figure 4. The method’s output is also compared against state-of-the-art low-light RGB image enhancement techniques to highlight the inherent advantages of using the IR spectrum in these conditions.
(a) Raw IR Input
(b) Denoised Output
(c) Ground-Truth
Figure 4: A qualitative comparison of the raw IR input with emitter patterns, the reconstructed image, and the corresponding ground-truth image.
As seen in the visual examples, the reconstructed output successfully removes the emitter patterns while preserving key features such as object boundaries and textures. The output image is visually coherent and appears cleaner, making it a much more suitable input for downstream computer vision algorithms. Even object shadows which are hard to see in the raw IR image can be seen clearly in the reconstructed image.
The advantages of the CLEAR-IR method are further emphasised through a quantitative comparison with traditional methods such as Contrast Limited Adaptive Histogram Equalisation (CLAHE) [27], Retinex Theory inspired techniques [28] and state-of-the-art low-light RGB image enhancement techniques such as Zero-Reference Deep Curve Estimation (Zero-DCE) [29], Color and Intensity Decoupling Network (CIDNet) [30], and LLFormer [31]. For this comparison, the similarity of each method’s output to a ground truth RGB image is evaluated using the Structural Similarity Index Measure (SSIM).
The visual results in Figure 5 shows the outputs from different techniques on the scene. While other methods attempt to enhance the image, the enhanced RGB images are affected by the effects of noise observed in low-light scenes. The IR cameras are more robust to the effect of this noise in low-light, as a result, the CLEAR-IR framework produces a clean, artifact-free image that maintains the scene’s structural integrity, as reflected by its superior SSIM score. This demonstrates that the proposed method is more effective at generating a representation of the scene suitable for downstream vision tasks.
TABLE I: Quantitative comparison of image enhancement methods against ground truth.
Method SSIM
Retinex Theory [27] 0.7418
CLAHE [28] 0.5672
Zero-DCE [29] 0.7624
LLFormer [30] 0.7063
CIDNet [31] 0.7788
CLEAR-IR (Proposed Method) 0.9521
The results, presented in Table I, show that the denoised IR output achieves a significantly higher SSIM score and is capable of maintaining the true structural integrity of the scene, even when compared to methods that produce a colour output.
Beyond the general quality metrics, the true utility of the proposed CLEAR-IR framework is demonstrated by its ability to facilitate high-level robotics tasks that are otherwise compromised by the structured light patterns and noise inherent in raw IR images and RGB images respectively. Table II presents a comprehensive quantitative comparison of the performance of various image enhancement methods on three high level tasks such as ORB feature matching, object detection, and ArUco marker detection. The results highlight a critical distinction between methods that simply improve image aesthetics and those that produce a robust, clean representation suitable for downstream algorithms.
(a) Ground Truth RGB Scene
(b) Dark RGB Scene
(c) CLAHE
(d) Retinex Theory
(e) Zero-DCE
(f) LLFormer
(g) CIDNet
(h) CLEAR-IR (Proposed Method)
Figure 5: Visual comparison of a ground truth RGB scene, a dark RGB scene, and the outputs of various image enhancement methods, including the proposed Denoised IR method.
TABLE II: Performance comparison of image enhancement methods for various computer vision tasks.
Method ORB Matches (Lowe’s Test) Object Detection (Visible Objects) ArUco Marker Detection
Correct Incorrect Detected Missed Mislabelled Correct Missed
Retinex Theory [27] 49 451 1 9 0 5 1
CLAHE [28] 36 464 1 9 0 3 3
Zero-DCE [29] 61 439 0 10 0 5 1
LLFormer [30] 53 447 0 10 0 5 1
CIDNet [31] 66 434 1 9 0 6 0
Raw IR 31 469 0 9 1 0 6
CLEAR-IR (Proposed Method) 75 342 7 2 1 6 0
IV-B Feature Matching and Keypoint Detection
To assess the impact of emitter pattern removal on traditional computer vision algorithms, the performance of the Oriented FAST and Rotated BRIEF (ORB) [32] feature detector was evaluated. ORB is a robust algorithm for finding keypoints and feature descriptors, which are fundamental for tasks such as Simultaneous Localisation and Mapping (SLAM) and visual odometry.
As shown in Figure 6, the raw IR images exhibit numerous false keypoint detections due to the high-frequency nature of the emitter pattern. These spurious features lead to inaccurate and unreliable feature matches, which would compromise downstream tasks. In contrast, the denoised images from the CLEAR-IR model provide a clean representation where ORB is able to detect and match meaningful keypoints on actual scene structures, such as object boundaries and corners. This demonstrates that the proposed method effectively suppresses noise that could be misinterpreted as texture by feature detectors, thereby producing a more reliable input for foundational vision algorithms.
(a) ORB matches on raw IR image
(b) ORB matches on denoised IR image
Figure 6: ORB feature matching on (a) raw IR and (b) denoised IR images.
IV-C Object Detection and Segmentation
To demonstrate the utility of the denoised images for advanced perception tasks, a pre-trained YOLOv8 object detection and segmentation model was applied. This model was not trained on IR data, making this a zero-shot evaluation to determine if the proposed preprocessing step can make the images usable with standard RGB-based vision pipelines.
(a) YOLOv8 on RGB image
(b) YOLOv8 on raw IR image
(c) YOLOv8 on Retinex image
(d) YOLOv8 on CLEAR-IR image
Figure 7: Object detection and segmentation results of YOLOv8 on raw and denoised IR images.
On the raw IR images, YOLOv8 failed to detect or segment objects. The emitter patterns and resulting image noise prevented the model from recognising the distinct shapes and features of common objects. After preprocessing the images with the CLEAR-IR model, the performance of YOLOv8 dramatically improved. As illustrated in Figure 7, the model successfully detected and segmented objects in the denoised images, which were completely missed in the raw inputs. While tested on a variety of images, some labels were incorrect since the YOLOv8 model had not been trained on the specific objects present in the scene. For example, traffic cones were occasionally misclassified as bottles, and a white storage cabinet was misclassified as a refrigerator. Similar misclassifications were also observed in the RGB stream, indicating that this is primarily a limitation of the YOLOv8 model rather than the proposed CLEAR-IR framework. Importantly, despite label mismatches, the model was able to consistently detect and track these objects across frames, demonstrating the usability of CLEAR-IR outputs for object-level perception.
IV-D ArUco Marker Detection
ArUco markers are fiducial markers widely used in robotics for localisation and pose estimation. The ability to reliably detect these markers is crucial for robust navigation. However, structured light patterns can interfere with the marker’s distinct black-and-white grid, often leading to detection failures. This interference is clearly visible in the raw infrared (IR) image shown in Figure 8b. For comparison, a baseline for successful detection is provided by the standard RGB image, as seen in Figure 8a.
(a) Markers in the RGB image
(b) No markers in the raw IR image due to structured light interference
(c) Markers in the CIDNet output
(d) Markers in the CLEAR-IR output
Figure 8: Comparison of ArUco marker detection performance on RGB, raw IR, and denoised IR images.
The CLEAR-IR framework addresses these challenges by enabling robust ArUco marker detection in IR images. By removing the spurious dot patterns, the model ensures the marker’s grid is clearly visible and interpretable by the detection algorithm, as demonstrated in Figure 8d. This is a significant advancement for low-light scenes where raw IR images are typically unusable for this task. The denoised output allows for consistent and accurate marker identification, thus enabling reliable localisation and mapping. This outcome directly contrasts with findings in prior work, such as [33], which concluded that raw IR images were unsuitable for low-light tag detection without the use of retroreflective materials. Similarly, this approach simplifies methods like [34], which propose using specialised tags for low-light localisation, by requiring no special tags or materials.
V Conclusion
This work demonstrates that the CLEAR-IR framework effectively addresses a fundamental challenge in robotic perception by successfully removing structured light emitter patterns from infrared images. Leveraging a U-Net-based architecture, the method produces visually coherent and artifact free outputs that are compatible with existing computer vision pipelines in a zero-shot scenario. The significance of this contribution is further evidenced by the substantial improvement in performance for crucial downstream tasks, including feature matching, object detection, and marker detection. As confirmed by quantitative results, this approach establishes a new standard, demonstrably surpassing state-of-the-art low-light RGB enhancement techniques. The CLEAR-IR framework thus provides a robust and reliable foundation for enabling autonomous systems to operate effectively in a wider range of challenging lighting conditions.
VI Future Direction
Future work will concentrate on two primary objectives to advance the CLEAR-IR framework’s utility in practical applications such as enhancing computational efficiency and enabling robust integration with visual Simultaneous Localisation and Mapping (vSLAM) systems. To improve processing time, research will explore model optimisation techniques, such as network quantisation and model pruning.
For integration with vSLAM, future work will focus on assessing the effectiveness of CLEAR-IR within a complete visual SLAM pipeline. This will involve real-time deployment and a quantitative evaluation of its impact on trajectory estimation accuracy and loop closure detection in illumination-constrained environments.
Acknowledgment
RAICo contributed to this work by providing funding and a supportive research environment. Jongyun Kim assisted the main author with setup and data collection in a dark environment.

### v2

@@
-CLEAR-IR: Clarity-Enhanced Active Reconstruction of Infrared Imagery
+CLEAR-IR: Clarity-Enhanced Active Reconstruction of Infrared Imagery for low-light enhancement
 Abstract
@@
 However, it is dominated by active emitter patterns that hinder high-level tasks such as object detection, tracking and localisation.
-To address this, a U-Net-based architecture is proposed that reconstructs clean IR images from emitter-populated input, improving both image quality and downstream robotic performance.
-This approach outperforms existing enhancement techniques and enables reliable operation of vision-driven robotic systems across illumination conditions from well-lit to extreme low-light scenes.
+To address this, a Deep Multi-scale Aware Overcomplete (DeepMAO) inspired architecture is proposed that reconstructs clean IR images from emitter populated input, improving both image quality and downstream robotic performance.
+This approach outperforms existing enhancement techniques and enables reliable operation of vision driven robotic systems across illumination conditions from well-lit to extreme low-light scenes.
+The results outline the ability of this work to be able to mimic RGB styling from the scene and its applicability on robotics tasks that were trained on RGB images, opening the possibility of doing these tasks in extreme low-light without on-board lighting.
 I Introduction
@@
 Attempts to compensate using an increased exposure time introduce motion blur, rendering them ineffective for dynamic environments.
-Furthermore, data-driven approaches that require training networks in long-exposure or low-light image pairs add complexity and reduce generalisability across different camera models.
-In addition, such methods are fundamentally limited by information-theoretic constraints and if critical scene information is not captured by the sensor due to noise or extreme darkness, no amount of training data can fully reconstruct it.
-The effectiveness of these networks is therefore heavily dependent on both data availability and inherent scene observability.
+Furthermore, data-driven approaches relying solely on visible spectrum low light and long exposure pairs are fundamentally limited by information-theoretic constraints.
+If critical scene information is not captured by the sensor due to extreme darkness or sensor level noise, it cannot be reconstructed regardless of the training data volume.
+Consequently, the efficacy of these networks remains strictly bound by the inherent observability of the visible spectrum in degraded conditions.
 Another approach involves equipping robots with onboard illumination systems [6].
@@
 As seen in Figure 1, the dark scene captured by the RGB camera fails to present good visual cues, whereas the same scene in the IR spectrum can be visualised with several objects in the scene.
-Although IR cameras are promising for dark-scene perception, they are not without their own limitations.
-Passive IR sensors, which operate without emitters, avoid structured-light artefacts but are typically expensive and limited to low-resolution outputs, making them impractical for robotics.
-Active IR systems, introduce structured light patterns that degrade image quality, interfere with feature extraction, and disrupt high-level perception tasks such as object detection, tracking, and SLAM.
-To overcome this, a deep learning-based framework, CLEAR-IR (Clarity-Enhanced Active Reconstruction of Infrared Imagery), is proposed.
-The framework is designed to reconstruct clean and interpretable IR images from raw emitter-populated inputs.
-The method leverages a U-Net architecture trained to suppress structured light noise while preserving spatial and textural details critical for vision-based tasks.
-The output is a denoised, visually coherent IR image that can serve as a robust input to standard perception modules.
-This is the first work to leverage the infrared stream for high-level robotic tasks by removing active emitter patterns in scenes ranging from well-illuminated to extreme low-light.
-The contributions of this work are:
-  • CLEAR-IR, a U-Net–based framework tailored for reconstructing clean IR images by suppressing structured-light artefacts, is introduced.
-  • A composite loss function is developed to preserve spatial and textural details critical for robotic perception while robustly removing high-frequency emitter patterns.
-  • Evaluation of IR image reconstruction for robotics is presented, demonstrating significant improvements across key tasks such as feature matching, object detection, and marker detection.
+Although IR cameras are less susceptible to such scattering due to its longer wavelength and are promising for dark-scene perception, they are not without their own limitations.
+Passive IR sensors, which operate without emitters, avoid structured-light artefacts but are typically expensive and limited to low-resolution outputs.
+Conversely, active IR systems often rely on structured light or flash patterns, and while these provide better penetration through dusty scenes than having an ambient light in extreme low-light scenes, the resulting high-intensity reflections and projected patterns can interfere with feature extraction, and disrupt downstream perception tasks such as object detection and Simultaneous Localisation and Mapping (SLAM).
 (a) Low-light room captured using an RGB camera
@@
 Figure 1: Low-light room captured using different cameras
+To overcome this, a deep learning-based framework, CLEAR-IR (Clarity-Enhanced Active Reconstruction of Infrared Imagery), is proposed.
+The framework is designed to reconstruct clean and interpretable IR images from raw emitter populated inputs.
+The method leverages a Deep Multi-scale Aware Overcomplete (DeepMAO [26]) inspired architecture trained to suppress structured light noise while preserving spatial and textural details critical for vision-based tasks.
+The output is a reconstructed, visually coherent IR image that can serve as a robust input to standard perception modules.
+This is the first work to leverage the infrared stream for high-level robotic tasks by removing active emitter patterns in scenes ranging from well-illuminated to extreme low-light.
+The contributions of this work are:
+  • CLEAR-IR, a DeepMAO inspired architecture tailored for reconstructing clean IR images by suppressing structured-light artefacts, is introduced.
+  • We utilise a composite loss function specifically selected to accommodate the spectral disparity and lack of pixel perfect alignment between IR and RGB inputs, allowing for robust pattern removal while preserving details critical for robotic perception.
+  • Evaluation of IR image reconstruction for robotics is presented, demonstrating significant improvements across key tasks such as object detection and marker detection.
+  • Comparison with other SOTA techniques for localisation in extreme low-light scenes.
 II Related Work
@@
 Thus, preserving fine-grained but genuine details is critical for ensuring robustness in downstream tasks such as vSLAM and visual odometry.
-Active infrared (IR) systems, which operate by illuminating a scene with an IR light source and capturing the reflected light, have been a viable alternative for dark-scene perception.
-These systems have been widely adopted in applications such as facial recognition [14] and depth sensing [15].
-However, the structured light patterns they use for scene illumination and depth estimation, while beneficial for some tasks, present a significant challenge for 2D vision tasks like object detection and segmentation.
-Repetitive high-frequency patterns can be misinterpreted as texture or objects by standard vision algorithms, leading to false positives or missed detections, which limits the amount of work done using these sensors.
-Previous works have only attempted to remove fixed patterns from IR images, as seen in [16, 17, 18].
-This just enhances the extremely noisy input of IR images, but does not focus on making it more useful for high-level robotics tasks.
-The U-Net architecture, with its symmetric encoder-decoder structure and skip connections, has shown remarkable success in a variety of image-to-image translation tasks, and was first introduced for biomedical image segmentation in [19].
-Its ability to capture both high-level semantic information from the encoder path and fine-grained spatial details from the decoder path makes it an ideal candidate for the specific denoising problem encountered in IR images.
+Active infrared (IR) systems provide a robust alternative for dark-scene perception by illuminating the environment with a dedicated source.
+While high power IR floodlights offer uniform coverage, they are often impractical for mobile robotics due to the inverse-square law, the power required for an adequate signal-to-noise ratio scales quadratically with distance.
+Furthermore, floodlights frequently cause shadows and also severe near-field saturation, where excessive photon flux annihilates local contrast and occludes critical features.
+To mitigate this while providing the structural texture necessary for navigation on featureless surfaces, depth camera manufacturers utilise structured-light or modulated laser patterns [1].
+Beyond their sensing advantages, these active IR systems are already standard components in many robotic platforms and leveraging them for 2D perception minimizes system complexity, reduces payload weight, and lowers maintenance costs compared to adding auxiliary illumination hardware.
+By distributing energy into discrete, evenly spaced points, these systems maintain a higher usable dynamic range even as pattern density increases at close proximity, the localised nature of the emission prevents the global washout characteristic of floodlights.
+However, these high-frequency patterns present a significant challenge for 2D vision tasks, as standard algorithms often misinterpret them as scene texture, leading to false detections.
+Previous denoising works [12, 21, 2] focus primarily on removing static sensor level artefacts rather than addressing the dynamic, pattern dependent noise inherent to these active emitters.
+Simply enhancing these inputs does not bridge the domain gap required for robotics, as static corrections cannot account for the characteristics discussed in Section III-A.
+The U-Net architecture, with its symmetric encoder-decoder structure, skip connections, has shown remarkable success in a variety of image-to-image translation tasks, and was first introduced for biomedical image segmentation in [24].
+Its ability to capture both high-level semantic information from the encoder path and fine-grained spatial details from the decoder path makes it an ideal candidate for the specific reconstruction problem encountered in IR images.
 This architecture has also been successfully adapted for other image reconstruction tasks, such as super-resolution [16] and general image denoising [34].
 In contrast to existing methods that primarily focus on low-light enhancement or general noise removal, this work addresses the specific and unique problem of structured light artefacts in active IR imagery.
-Using the U-Net architecture, it is possible to effectively learn to suppress these specific patterns while preserving the crucial underlying semantic information for high-level vision tasks.
+Using the DeepMAO architecture, it is possible to effectively learn to suppress these specific patterns while preserving the crucial underlying semantic information for high-level vision tasks.
 III Methodology
 This section presents the CLEAR-IR framework.
-This method leverages a U-Net architecture and a composite loss function to effectively remove structured light artefacts while preserving critical visual details.
+Unlike standard encoder-decoder approaches, CLEAR-IR uses a hybrid dual-branch architecture to effectively remove structured light artefacts while preserving critical visual details.
 III-A Emitter Pattern Characteristics
@@
 Figure 2: The key factors influencing the appearance of the structured light emitter pattern in active infrared imagery.
-This variability in pattern appearance across different scenes creates a significant challenge in selecting adequate losses that can remove them from the images while being scene-agnostic.
-III-B U-Net Architecture for Emitter Pattern Removal
-To effectively reconstruct the image and remove the noise caused by the emitter pattern, a standard U-Net architecture was used.
-This convolutional neural network is particularly well suited for image-to-image translation tasks due to its symmetric encoder-decoder structure and strategic use of skip connections, as depicted in Figure 3.
-The network is configured to process greyscale images with an input resolution of 640x480 pixels.
-The encoder path progressively downsamples the image, using a series of convolutional and max-pooling layers to extract increasingly abstract features and high-level contextual information.
-The decoder path then reconstructs the image using transposed convolutional layers to upsample the features back to the original resolution.
-A crucial element of this architecture are its skip connections, which concatenate the feature maps from the encoder to the corresponding layers in the decoder.
-These connections are vital for preserving fine-grained spatial details that might otherwise be lost during the downsampling process.
-The network combines these features at each upsampling stage to produce a final, single-channel output image that is the same size as the input, with the emitter pattern effectively removed.
-Figure 3: The U-Net architecture used for image reconstruction and emitter pattern removal.
-III-C Dataset and Data Augmentation
-A custom paired dataset of IR images and their corresponding clean ground-truth RGB images converted to greyscale was created for model training using an Intel Realsense D455 camera [22].
-Data collection was performed in an indoor workspace with varied lighting conditions.
-The illumination levels and object distance were intentionally adjusted to capture emitter patterns with different sharpness and intensity, creating a varied dataset.
-The initial dataset contained 10,000 paired images.
-To enhance model robustness and prevent overfitting, a data augmentation process was applied.
-For each image pair, five augmented versions were generated using random transformations, which included:
-  • Rotation: Up to 30 degrees in either direction.
-  • Flipping: Random horizontal and vertical flips.
-  • Affine Transformations: Minor rotations and translations.
-  • Zoom: Random zooming in by up to 30%.
-  • Brightness and Contrast: Adjustments to simulate lighting variations.
-The final data had 50000 paired sets of IR and greyscale RGB images and will be released as an open-sourced dataset.
-III-D Composite Loss Function and Training
-To produce high-quality image reconstructions, a composite loss function was used to guide the network’s training.
-This function is a weighted sum of multiple components, each designed to optimise a different aspect of the output, ensuring that the reconstructed image is not pixel accurate but perceptually and structurally similar to the ground truth.
-The main reason for choosing these losses besides mean absolute or mean squared error was due to the epipolar transformation of pixels between the IR and RGB images.
-III-D1 Loss Components
-The total loss is a weighted sum of several components.
-Let $\alpha,\beta,\gamma,\delta,\epsilon$, and $\zeta$ be the empirically tuned weighting coefficients for each corresponding loss component.
-The total loss is defined as:
-[…]
-  • Mean Absolute Error (MAE): This is a pixel-wise loss that minimises the average absolute difference between the predicted image $\hat{I}$ and the ground truth image $I$.
-It provides a foundational measure of pixel level accuracy.
-    $\mathcal{L}_{mae}=\frac{1}{N}\sum_{i=1}^{N}|\hat{I}_{i}-I_{i}|$ (4)
-  • Structural Similarity Index (SSIM) Loss: This loss measures the similarity in luminance, contrast, and structure between images.
-By minimising this loss, the network is encouraged to preserve the overall structural integrity of the scene [23].
-    $\mathcal{L}_{ssim}=1-\text{SSIM}(\hat{I},I)$ (5)
-  • High-Frequency and Sobel Edge Losses: These losses are used to preserve fine details and sharp boundaries.
-The high-frequency loss uses a Laplacian filter to penalise differences in the high-frequency components of the images, while the Sobel edge loss penalises differences in image gradients to ensure edges are accurately reproduced.
-    $\mathcal{L}_{freq}=\left\|(\hat{I}-I)*K_{Laplacian}\right\|_{1}$ (6)
-    $\mathcal{L}_{sobel}=\left\|\nabla_{x}\hat{I}-\nabla_{x}I\right\|_{1}+\left\|\nabla_{y}\hat{I}-\nabla_{y}I\right\|_{1}$ (7)
-  • Perceptual Loss: This loss component uses a pre-trained VGG19 network as a feature extractor.
-The network’s weights, which were pre-trained on the ImageNet dataset [24], are frozen during training.
-This allows the model to leverage learned feature representations that are more aligned with human visual perception than simple pixel values.
-To apply this to greyscale images, the input is converted to a 3 channel format by concatenating the single channel three times.
-Instead of comparing pixels directly, this loss compares the feature maps from a set of specific intermediate layers within the VGG19 network.
-By using layers such as block3_conv2, block4_conv2, and the layers within block5, the loss encourages the reconstructed image to have similar high level content and style as the ground truth.
-This is crucial for generating an output that is visually convincing and suitable for high level vision tasks [25].
-    $\mathcal{L}_{perceptual}=\sum_{j}\|\Phi_{j}(\hat{I})-\Phi_{j}(I)\|_{1}$ (8)
-    where $\Phi_{j}(\cdot)$ is the feature map from layer $j$ of the VGG19 network.
-  • Total Variation (TV) Loss: This regularizer promotes spatial smoothness and helps to reduce spurious noise in the output image [26].
-    $\mathcal{L}_{tv}=\sum_{i,j}\left(\left(\hat{I}_{i+1,j}-\hat{I}_{i,j}\right)^{2}+\left(\hat{I}_{i,j+1}-\hat{I}_{i,j}\right)^{2}\right)$ (9)
-III-D2 Weighting Strategy
-The weighting coefficients for each component ($\alpha,\beta,\gamma,\delta,\epsilon,\zeta$) were experimentally tuned to balance their influence on the final output.
-The primary goal was to prioritise visual quality and structural integrity over a strict pixel-by-pixel match.
-High weights were assigned to $\mathcal{L}_{ssim}$ to preserve scene structure, while a significant but not overwhelming weight was given to $\mathcal{L}_{perceptual}$ to ensure the output was visually convincing.
-Moderate weights were used for $\mathcal{L}_{freq}$ and $\mathcal{L}_{sobel}$ to preserve fine details and edges while also not prioritising emitter patterns which can be classed as high-frequency patterns in the image.
-In contrast, very low weights were given to $\mathcal{L}_{mae}$ and $\mathcal{L}_{tv}$, as their purpose was to act as subtle regularizers rather than primary training objectives.
-III-E Preprocessing and Training
-After augmentation, the dataset, which contained 50,000 paired images, 90% of the data was split into a training set and the remaining 10% into the validation set.
-All images were resized to a consistent resolution of 640x480 pixels and normalised to the range of [0, 1].
-The model was trained using the Adam optimizer with a CosineDecayRestarts learning rate schedule to minimise the composite loss function.
-To prevent overfitting, an early stopping mechanism with a patience of 30 epochs was implemented.
-The model was trained for a total of 126 epochs, at which point it met the early stopping criteria, indicating that a sufficient level of convergence had been reached.
-IV Experimental Results
-This section presents a comprehensive evaluation of the proposed CLEAR-IR framework, focusing on its effectiveness in preparing infrared images for high-level computer vision tasks.
-Since this is the first work to specifically address the removal of active emitter patterns, the primary evaluation is a qualitative and quantitative comparison of the performance of enhancement techniques on the RGB stream versus the denoised images of the IR stream produced by this work.
-IV-A Qualitative and Quantitative Analysis
-To provide a visual assessment of the method’s reconstruction quality and its robustness in low light conditions, a qualitative comparison is first presented of a raw IR input, the model’s reconstructed output, and the corresponding ground truth image, as seen in Figure 4.
-The method’s output is also compared against state-of-the-art low-light RGB image enhancement techniques to highlight the inherent advantages of using the IR spectrum in these conditions.
-(a) Raw IR Input
-(b) Denoised Output
-(c) Ground-Truth
-Figure 4: A qualitative comparison of the raw IR input with emitter patterns, the reconstructed image, and the corresponding ground-truth image.
-As seen in the visual examples, the reconstructed output successfully removes the emitter patterns while preserving key features such as object boundaries and textures.
-The output image is visually coherent and appears cleaner, making it a much more suitable input for downstream computer vision algorithms.
-Even object shadows which are hard to see in the raw IR image can be seen clearly in the reconstructed image.
-The advantages of the CLEAR-IR method are further emphasised through a quantitative comparison with traditional methods such as Contrast Limited Adaptive Histogram Equalisation (CLAHE) [27], Retinex Theory inspired techniques [28] and state-of-the-art low-light RGB image enhancement techniques such as Zero-Reference Deep Curve Estimation (Zero-DCE) [29], Color and Intensity Decoupling Network (CIDNet) [30], and LLFormer [31].
-For this comparison, the similarity of each method’s output to a ground truth RGB image is evaluated using the Structural Similarity Index Measure (SSIM).
-The visual results in Figure 5 shows the outputs from different techniques on the scene.
-While other methods attempt to enhance the image, the enhanced RGB images are affected by the effects of noise observed in low-light scenes.
-The IR cameras are more robust to the effect of this noise in low-light, as a result, the CLEAR-IR framework produces a clean, artifact-free image that maintains the scene’s structural integrity, as reflected by its superior SSIM score.
-This demonstrates that the proposed method is more effective at generating a representation of the scene suitable for downstream vision tasks.
-TABLE I: Quantitative comparison of image enhancement methods against ground truth.
-Method SSIM
-Retinex Theory [27] 0.7418
-CLAHE [28] 0.5672
-Zero-DCE [29] 0.7624
-LLFormer [30] 0.7063
-CIDNet [31] 0.7788
-CLEAR-IR (Proposed Method) 0.9521
-The results, presented in Table I, show that the denoised IR output achieves a significantly higher SSIM score and is capable of maintaining the true structural integrity of the scene, even when compared to methods that produce a colour output.
-Beyond the general quality metrics, the true utility of the proposed CLEAR-IR framework is demonstrated by its ability to facilitate high-level robotics tasks that are otherwise compromised by the structured light patterns and noise inherent in raw IR images and RGB images respectively.
-Table II presents a comprehensive quantitative comparison of the performance of various image enhancement methods on three high level tasks such as ORB feature matching, object detection, and ArUco marker detection.
-The results highlight a critical distinction between methods that simply improve image aesthetics and those that produce a robust, clean representation suitable for downstream algorithms.
+This variability in pattern appearance across different scenes poses a challenge in selecting adequate losses that can remove them from the images while being scene agnostic.
+III-B CLEAR-IR Architecture
+To preserve high-frequency spatial details, this work adopts a modified DeepMAO architecture.
+Although the original DeepMAO was modelled for satellite image segmentation using a pretrained EfficientNet backbpne along with a combination of focal loss and dice loss, this work adapts its dual-stream ideology for IR image reconstruction.
+This work simplifies the original architecture by replacing the heavy encoder with a U-Net backbone and we simplify the multi-scale overcomplete (MSO) module from non-integer interpolations to a sequence of full-resolution convolution layers.
+We retain the original architecture’s additive residual fusion to facilitate a direct gradient flow for fine-grained textures.
+Furthermore, we transition the training objective from pixel-level classification to intensity regression, utilising a combined perceptual and structure loss tailored for scene reconstruction.
+III-B1 Context Stream (U-Net Backbone)
+Unlike the original DeepMAO which relied on a heavy EfficientNet-B3 encoder, our first branch utilises a U-Net backbone designed to efficiently capture global context and low-frequency structural information.
+  • Hierarchical Encoding: The encoder processes the input through a series of convolutional blocks with increasing filter depths (32, 64, 128) and $2\times 2$ max-pooling layers.
+This effectively expands the receptive field, allowing the network to understand global scene geometry.
+  • Regularisation: To ensure stability during training and prevent overfitting on specific emitter patterns, dropout with a rate of $0.2$ is integrated into the deeper layers of the encoder and bottleneck.
+  • Decoder and Skip Connections: The decoder utilises transposed convolutions to upsample feature maps, concatenated with corresponding features from the encoder to recover spatial information lost during pooling.
 (a) Ground Truth RGB Scene
@@
 (g) CIDNet
-(h) CLEAR-IR (Proposed Method)
-Figure 5: Visual comparison of a ground truth RGB scene, a dark RGB scene, and the outputs of various image enhancement methods, including the proposed Denoised IR method.
-TABLE II: Performance comparison of image enhancement methods for various computer vision tasks.
-Method ORB Matches (Lowe’s Test) Object Detection (Visible Objects) ArUco Marker Detection
-Correct Incorrect Detected Missed Mislabelled Correct Missed
-Retinex Theory [27] 49 451 1 9 0 5 1
-CLAHE [28] 36 464 1 9 0 3 3
-Zero-DCE [29] 61 439 0 10 0 5 1
-LLFormer [30] 53 447 0 10 0 5 1
-CIDNet [31] 66 434 1 9 0 6 0
-Raw IR 31 469 0 9 1 0 6
-CLEAR-IR (Proposed Method) 75 342 7 2 1 6 0
-IV-B Feature Matching and Keypoint Detection
-To assess the impact of emitter pattern removal on traditional computer vision algorithms, the performance of the Oriented FAST and Rotated BRIEF (ORB) [32] feature detector was evaluated.
-ORB is a robust algorithm for finding keypoints and feature descriptors, which are fundamental for tasks such as Simultaneous Localisation and Mapping (SLAM) and visual odometry.
-As shown in Figure 6, the raw IR images exhibit numerous false keypoint detections due to the high-frequency nature of the emitter pattern.
-These spurious features lead to inaccurate and unreliable feature matches, which would compromise downstream tasks.
-In contrast, the denoised images from the CLEAR-IR model provide a clean representation where ORB is able to detect and match meaningful keypoints on actual scene structures, such as object boundaries and corners.
-This demonstrates that the proposed method effectively suppresses noise that could be misinterpreted as texture by feature detectors, thereby producing a more reliable input for foundational vision algorithms.
-(a) ORB matches on raw IR image
-(b) ORB matches on denoised IR image
-Figure 6: ORB feature matching on (a) raw IR and (b) denoised IR images.
-IV-C Object Detection and Segmentation
-To demonstrate the utility of the denoised images for advanced perception tasks, a pre-trained YOLOv8 object detection and segmentation model was applied.
+(h) Raw IR
+(i) U-Net
+(j) CLEAR-IR (Proposed Method)
+Figure 3: Visual comparison of enhancement methods on example scene.
+III-B2 Detail Stream (Overcomplete Branch)
+While the pooling operations in the U-Net are effective at smoothing out the artificial structured light patterns, they are also destructive to the true high frequency details of the scene.
+To mitigate this, the second branch acts as an overcomplete pathway.
+  • Full Spatial Resolution: In the original DeepMAO, the overcomplete path utilized complex non-integer interpolations and dilated convolutions.
+We simplify this mechanism to better suit image reconstruction.
+Our branch accepts the initial feature maps (32 filters) and processes them through a cascade of convolutional layers without any scaling, pooling, or strided operations.
+  • Detail Preservation and Noise Suppression: By strictly maintaining a 1:1 input-to-output spatial correspondence, this branch bypasses the destructive downsampling of the U-Net.
+It specialises in isolating the true high-frequency structural boundaries from the structured light noise, ensuring that sharp edges and fine scene textures are preserved and injected back into the final reconstruction.
+III-B3 Feature Fusion
+The outputs of the U-Net branch and the spatial overcomplete branch are integrated using an additive fusion strategy, mirroring the residual nature of the original network.
+$F_{fused}=F_{unet}\oplus F_{overcomplete}$ (1)
+where $\oplus$ denotes element-wise addition.
+This residual-style fusion allows the network to combine the clean structural predictions from the U-Net with the texture corrections from the overcomplete branch.
+The fused features are passed through a final convolutional refinement block to generate the continuous intensity values of the reconstructed IR image.
+III-C Loss Function
+Despite the extrinsic calibration and warping process, minor alignment residuals persist due to parallax and sensor baseline between the IR and RGB camera.
+Therefore, this work uses a composite loss function that prioritises perceptual quality and structural integrity over simple pixel-wise accuracy.
+The total loss $\mathcal{L}_{\text{total}}$ is a weighted sum of six components:
+[…]
+The components are defined as follows:
+  • Spatial Fidelity ($\mathcal{L}_{mae},\mathcal{L}_{ssim}$): This work use Mean Absolute Error (MAE) for base intensity matching and Structural Similarity Index (SSIM) [30] to preserve luminance and contrast structure.
+  • Detail Preservation ($\mathcal{L}_{freq},\mathcal{L}_{sobel}$): To retain sharp edges, $\mathcal{L}_{freq}$ minimizes the $L_{1}$ distance between Laplacian-filtered images, while $\mathcal{L}_{sobel}$ penalizes discrepancies in image gradients.
+  • Perceptual Quality ($\mathcal{L}_{perceptual}$): A VGG19 network pre-trained on ImageNet [7] is used as a feature extractor.
+The loss minimises the $L_{1}$ distance between feature maps of specific intermediate layers, ensuring the reconstruction aligns with human visual perception [18]:
+  • Regularization ($\mathcal{L}_{tv}$): Total Variation (TV) loss promotes spatial smoothness to reduce spurious noise in uniform regions.
+The weighting coefficients ($\alpha,\beta,\gamma,\delta,\epsilon,\zeta$) were experimentally tuned to prioritize $\mathcal{L}_{ssim}$ and $\mathcal{L}_{perceptual}$, ensuring the model focuses on structural removal of the pattern rather than overfitting to minor pixel misalignments.
+III-D Data Augmentation and Training Implementation
+A custom dataset was developed by capturing 6,719 initial pairs of active IR and greyscale RGB images using an Intel RealSense D455 camera [17].
+To ensure spatial correspondence, we utilised factory calibrated extrinsics to align the sensors.
+These base frames were diversified through data augmentation, including random rotations ($\pm 30^{\circ}$), flipping, affine transformations, and brightness adjustments resulting in a final dataset of 33,595 images.
+The model was trained to map IR inputs to an RGB-style ground truth, enabling the deployment of downstream vision tasks originally developed for the visible spectrum.
+The network was trained on a 90 to 10 data split at $640\times 480$ resolution using the Adam optimizer with a cosine decay scheduler.
+Early stopping was also used to ensure convergence while preventing overfitting.
+IV Experimental Results
+This section presents a comprehensive evaluation of the proposed CLEAR-IR framework, focusing on its effectiveness in preparing infrared images for high-level computer vision tasks.
+Since this is the first work to specifically address the removal of active emitter patterns, the primary evaluation is a qualitative and quantitative comparison of the performance of enhancement techniques on the RGB stream versus the reconstructed images of the IR stream produced by this work.
+The advantages of the CLEAR-IR method is emphasised through a quantitative comparison against both traditional baselines and a diverse set of state of the art (SOTA) enhancement methods.
+We evaluate against classical spatial-domain techniques, including CLAHE [27] and retinex theory inspired methods [3].
+To benchmark against modern deep learning advances, this work evaluates Zero-DCE [13], a zero-reference method for unsupervised curve estimation, LLFormer [29], a benchmark Transformer-based architecture optimised for long-range dependency modelling and CIDNet [31], a multi-axis decoupling network that utilises a new HVI colour space to suppress colour bias and noise.
+These methods provide a robust evaluation across the leading unsupervised, attention-based, and colour-decoupling methodologies.
+The visual results in Figure 3 shows the outputs from different techniques on an unknown scene.
+While other methods attempt to enhance the image, the enhanced RGB images are affected by the effects of noise observed in low-light scenes.
+The IR cameras are more robust to the effect of this noise in low-light, as a result, the CLEAR-IR framework produces a clean, artifact free image that maintains the scene’s structural integrity without introducing any additional artefacts or hallucinating.
+This demonstrates that the proposed method is more effective at generating a representation of the scene suitable for downstream vision tasks.
+TABLE I: Performance comparison of image enhancement methods for various computer vision tasks.
+Method Object Detection (Visible Objects) ArUco Marker Detection
+Detected Missed Mislabelled Correct Missed
+Retinex Theory [3] 1 9 0 5 1
+CLAHE [27] 1 9 0 3 3
+Zero-DCE [13] 0 10 0 5 1
+LLFormer [29] 0 10 0 5 1
+CIDNet [31] 1 9 0 6 0
+Raw IR 0 9 1 0 6
+U-Net 7 3 0 6 0
+CLEAR-IR (Proposed Method) 11 0 2 6 0
+Beyond the general quality metrics, the true utility of the proposed CLEAR-IR framework is demonstrated by its ability to facilitate high-level robotics tasks that are otherwise compromised by the structured light patterns and noise inherent in raw IR images and RGB images respectively.
+Table I presents a comprehensive quantitative comparison of the performance of various image enhancement methods on high level tasks such as object detection, and ArUco marker detection.
+The results highlight a critical distinction between methods that simply improve image aesthetics and those that produce a robust, clean representation suitable for downstream algorithms.
+The proposed work is also used in feature tracking and odometry estimation to highlight the stability across frames in unseen environments.
+IV-A Object Detection and Segmentation
+To demonstrate the utility of the reconstructed images for advanced perception tasks, a pre-trained YOLOv26 object detection and segmentation model was used for evaluation.
 This model was not trained on IR data, making this a zero-shot evaluation to determine if the proposed preprocessing step can make the images usable with standard RGB-based vision pipelines.
-(a) YOLOv8 on RGB image
-(b) YOLOv8 on raw IR image
-(c) YOLOv8 on Retinex image
-(d) YOLOv8 on CLEAR-IR image
-Figure 7: Object detection and segmentation results of YOLOv8 on raw and denoised IR images.
-On the raw IR images, YOLOv8 failed to detect or segment objects.
+(a) YOLOv26 on RGB image
+(b) YOLOv26 on raw IR image
+(c) YOLOv26 on U-Net enhanced image
+(d) YOLOv26 on CLEAR-IR image
+Figure 4: Object detection and segmentation results of YOLOv26 on raw and reconstructed IR images.
+On the raw IR images, YOLOv26 failed to detect or segment objects.
 The emitter patterns and resulting image noise prevented the model from recognising the distinct shapes and features of common objects.
-After preprocessing the images with the CLEAR-IR model, the performance of YOLOv8 dramatically improved.
-As illustrated in Figure 7, the model successfully detected and segmented objects in the denoised images, which were completely missed in the raw inputs.
-While tested on a variety of images, some labels were incorrect since the YOLOv8 model had not been trained on the specific objects present in the scene.
+After preprocessing the images with the CLEAR-IR model, the performance of YOLOv26 dramatically improved.
+As illustrated in Figure 4, the model successfully detected and segmented objects in the reconstructed images, which were completely missed in the raw inputs.
+While tested on a variety of images, some labels were incorrect since the YOLOv26 model had not been trained on the specific objects present in the scene.
 For example, traffic cones were occasionally misclassified as bottles, and a white storage cabinet was misclassified as a refrigerator.
-Similar misclassifications were also observed in the RGB stream, indicating that this is primarily a limitation of the YOLOv8 model rather than the proposed CLEAR-IR framework.
-Importantly, despite label mismatches, the model was able to consistently detect and track these objects across frames, demonstrating the usability of CLEAR-IR outputs for object-level perception.
-IV-D ArUco Marker Detection
+Although the greyscale nature of IR data makes some misclassification inevitable without fine-tuning.
+Importantly, despite some label mismatches, the model was able to consistently detect and track these objects across frames, demonstrating the usability of CLEAR-IR outputs for object-level perception.
+IV-B ArUco Marker Detection
 ArUco markers are fiducial markers widely used in robotics for localisation and pose estimation.
@@
 (d) Markers in the CLEAR-IR output
-Figure 8: Comparison of ArUco marker detection performance on RGB, raw IR, and denoised IR images.
+Figure 5: Comparison of ArUco marker detection performance on RGB, raw IR, and reconstructed IR images.
 The CLEAR-IR framework addresses these challenges by enabling robust ArUco marker detection in IR images.
@@
 This is a significant advancement for low-light scenes where raw IR images are typically unusable for this task.
-The denoised output allows for consistent and accurate marker identification, thus enabling reliable localisation and mapping.
+The reconstructed output allows for consistent and accurate marker identification, thus enabling reliable localisation and mapping.
 This outcome directly contrasts with findings in prior work, such as [9], which concluded that raw IR images were unsuitable for low-light tag detection without the use of retroreflective materials.
 Similarly, this approach simplifies methods like [25], which propose using specialised tags for low-light localisation, by requiring no special tags or materials.
+IV-C Visual Simultaneous Localisation and Mapping (VSLAM) Performance
+TABLE II: Quantitative evaluation of SLAM trajectory accuracy (RMSE in meters) and inference latency across Low Light (LL) and Extreme Low Light (ELL) scenarios.
+Method Low Light (LL) - RMSE $\pm$ std Extreme Low Light (ELL) - RMSE $\pm$ std Inference (ms)
+FB SQ1 SQ2 FB SQ1 SQ2 Mean std
+[…]
+[…]
+[…]
+[…]
+[…]
+Raw IR DNI DNI DNI DNI DNI DNI DNI DNI
+[…]
+[…]
+DNI: Did Not Initialise; Bold shows best results
+(a) SQ1 trajectories in the low-light sequence
+(b) SQ1 trajectories in the extreme low-light sequence
+Figure 6: Square 1 lap (SQ1) Trajectories
+To evaluate the practical utility of CLEAR-IR for robotics applications, we integrated the reconstructed infrared stream into a modified ORB-SLAM3 [4] Visual SLAM pipeline.
+In our implementation, the default ORB feature extractor was replaced with SuperPoint [8], and feature matching was adapted accordingly to use floating-point descriptor matching with L2 distance.
+This modification ensures compatibility with learned feature descriptors while preserving the full SLAM pipeline, including tracking, local mapping, and loop closure.
+Robust SLAM performance critically depends on stable and repeatable feature tracking across frames.
+In raw IR images, however, the structured light emitter pattern introduces dense artificial high-frequency responses.
+These phantom features are inconsistent across viewpoints and frames, leading to incorrect correspondences, tracking divergence, and eventual tracking failure.
+This behaviour is reflected in the Table II where Raw IR fails in all sequences.
+To ensure fair comparison with RGB-based enhancement methods, monocular RGB and depth sequences were recaptured under controlled lighting conditions. Three trajectories were recorded:
+  • Forward-Backward (FB): This sequence is a simple movement of moving in a straight line till an object and reversing backward
+  • Square 1 lap (SQ1): This sequence is a simple movement of rotating along an arena and returning to the start point as seen in Figure 6.
+  • Square 2 lap (SQ2): This sequence repeats the SQ1 trajectory for two full laps, providing a benchmark for observing the model’s ability to facilitate successful loop closure and drift correction
+Table II reports results in low-light conditions where limited ambient illumination is still present.
+Under these conditions, the RGB enhancement methods (Retinex, CLAHE, Zero-DCE, LLFormer, CIDNet) perform reliably because the RGB images retain sufficient structural information for feature extraction and tracking.
+CLEAR-IR and the U-Net baseline achieve performance comparable to these RGB-based methods.
+In simpler trajectories (FB and SQ1), CLEAR-IR performs on par with the RGB enhancement models.
+Notably, in the SQ2 sequence, CLEAR-IR achieves the lowest RMSE among all evaluated methods.
+Since SQ2 stresses loop closure and long-term consistency, this result indicates that CLEAR-IR reconstructs temporally stable structures that reduce drift accumulation over extended trajectories.
+Importantly, unlike raw IR, the reconstructed outputs do not introduce false high-frequency artefacts from the emitter pattern.
+Instead, the learned reconstruction produces geometrically meaningful and temporally consistent features suitable for SLAM.
+Furthermore in extreme low-light conditions the RGB images become severely underexposed, containing insufficient texture or gradient information for reliable feature detection.
+Consequently, all RGB-based enhancement methods fail to maintain tracking across every sequence, as indicated by missing RMSE values.
+In contrast, the IR images remain unaffected by ambient lighting conditions.
+While raw IR still fails due to emitter induced phantom features, both U-Net and CLEAR-IR successfully enable full SLAM operation across all sequences.
+CLEAR-IR consistently outperforms the U-Net baseline in terms of RMSE, achieving the best accuracy in FB, SQ1, and SQ2.
+This demonstrates a key advantage of the proposed framework: when RGB information collapses, CLEAR-IR leverages the invariant structural information present in the IR modality to reconstruct geometrically consistent representations that remain usable for feature-based SLAM.
+In terms of inference latency, CLEAR-IR maintains competitive runtime performance.
+While classical methods such as CLAHE are significantly faster, they do not provide robustness under extreme lighting degradation.
+Conversely, transformer-based models such as LLFormer incur substantially higher computational cost.
+CLEAR-IR achieves a favourable trade-off between localisation accuracy and inference speed, maintaining compatibility with real-time SLAM update requirements.
 V Conclusion
-This work demonstrates that the CLEAR-IR framework effectively addresses a fundamental challenge in robotic perception by successfully removing structured light emitter patterns from infrared images.
-Leveraging a U-Net-based architecture, the method produces visually coherent and artifact free outputs that are compatible with existing computer vision pipelines in a zero-shot scenario.
-The significance of this contribution is further evidenced by the substantial improvement in performance for crucial downstream tasks, including feature matching, object detection, and marker detection.
-As confirmed by quantitative results, this approach establishes a new standard, demonstrably surpassing state-of-the-art low-light RGB enhancement techniques.
+Utilising the infrared cameras already commonly integrated into robotic systems, the CLEAR-IR framework effectively addresses a fundamental perception challenge by removing emitter patterns from infrared images, successfully enabling low-light vision and extending the use case of these camera systems.
+Leveraging a DeepMAO inspired architecture, the method produces visually coherent and artifact free outputs that are compatible with existing computer vision pipelines in a fully dark environment.
+The significance of this contribution is further evidenced by the substantial improvement in performance for downstream tasks, including VIO, object detection, and marker detection.
+As confirmed by quantitative results, this approach establishes a new standard, demonstrably surpassing state-of-the-art low-light RGB enhancement techniques, showing flexibility even in entirely dark scenes.
 The CLEAR-IR framework thus provides a robust and reliable foundation for enabling autonomous systems to operate effectively in a wider range of challenging lighting conditions.
@@
 To improve processing time, research will explore model optimisation techniques, such as network quantisation and model pruning.
-For integration with vSLAM, future work will focus on assessing the effectiveness of CLEAR-IR within a complete visual SLAM pipeline.
-This will involve real-time deployment and a quantitative evaluation of its impact on trajectory estimation accuracy and loop closure detection in illumination-constrained environments.
 Acknowledgment
-RAICo contributed to this work by providing funding and a supportive research environment. Jongyun Kim assisted the main author with setup and data collection in a dark environment.
+RAICo contributed to this work by providing funding and a supportive research environment. Jongyun Kim assisted the authors with setup and data collection in a dark environment.